Suppose approximately 75% of all marketing personnel are extroverts, whereas about 50% of all computer programmers are introverts. (For each answer, enter a number. Round your answers to three decimal places.)

What is the probability that 2 or more are introverts?


4 computer programers
Answer #1

probability that 2 or more are introverts =P(X>=2) =1-P(X<=1)=1-P(X=0)-P(X=1)

=1-4C0(0.5)0(0.5)4-4C1(0.5)1(0.5)3=0.688